OpenStudy (anonymous):

Heather's family bought tickets at a local carnival. Her brother took1/5 the number of tickets that her cousin took, which was half the number Heather took. If her brother took 10 tickets, how many tickets did Heather take?

OpenStudy (anonymous):

All right. We know that Heather's brother took ten tickets. Heather's brother took (1/5) the number of tickets that her cousin took. So, we know that the number of tickets the cousin took is: 10 = x * (1/5)

OpenStudy (anonymous):

Once we solve for x. We know that x is half the number that Heather took. x * 2 = y Where why is how many tickets Heather took.

OpenStudy (anonymous):

im sill confused

OpenStudy (anonymous):

Okay. Heather's brother took ten tickets, which is (1/5) the number of tickets that the cousin took. We can also say that her cousin took 5 times the number of tickets as her brother... x = 10 * 5 is the same as 10 = x * 1/5

OpenStudy (anonymous):

So, how many tickets did her cousin take?

OpenStudy (anonymous):

half

OpenStudy (anonymous):

Heather took half the tickets that her cousin took. So, let's not worry about the half yet. Lets focus on this part: Her brother took1/5 the number of tickets that her cousin took

OpenStudy (anonymous):

Brother took 10. So 10 = (1/5) the number the cousin took.

OpenStudy (anonymous):

ooh thankk you

OpenStudy (anonymous):

How many did the cousin take?

OpenStudy (anonymous):

10

OpenStudy (anonymous):

The brother took ten...

OpenStudy (anonymous):

Brother = 10 Cousin = 10 * 5 Heather = Cousin * 2

OpenStudy (anonymous):

so its 50

OpenStudy (anonymous):

The cousin has 50, yes. How many does heather have?

OpenStudy (anonymous):

10